OpenStudy (anonymous):

URGENT!! for the quadratic equation it says,,, The solutions of ax^2 +bx+c=0,where a (then a funny sign!) 0, are given by...(then it says the formula)..... what is that funny sign??? and what if the sum is in minus?? and last but not least wat does that +- sign mean?? THNX!!

OpenStudy (whpalmer4):

\[ax^2+bx+c=0, a\ne0\]Is that what you see? the funny sign means "does not equal"

OpenStudy (whpalmer4):

the solutions are given by \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

ye after the a is the funny sign

OpenStudy (whpalmer4):

the \(\pm\) means "plus or minus" — there are actually two solutions, one with the plus, one with the minus.

OpenStudy (whpalmer4):

It is just a more compact way to write it than \[x = \frac{-b + \sqrt{b^2-4ac}}{2a},x=\frac{-b-\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

oh,thnx so we shud give two answers?? and work it out twice? and what if the sum your given is in minus is it the same method...

OpenStudy (whpalmer4):

Yes, a quadratic always has 2 solutions (though in some cases they are the same value). I'm not sure what you mean by "if the sum your given is in minus" — can you give me an example?

OpenStudy (anonymous):

sure.lets say\[3x^2-4x-2=0\]

OpenStudy (whpalmer4):

Okay, in this case, \(a = 3, b= -4, c = -2\)

OpenStudy (anonymous):

oh im so stupid.obviously so then you work out wats in the square root,then square root it,then -b plus the answer,then share by 2a?

OpenStudy (whpalmer4):

For your example: \[x = \frac{-(-4)\pm\sqrt{(-4)^2-4(3)(-2)}}{2*3} = \frac{4\pm\sqrt{16+24}}{6}=\frac{4\pm2\sqrt{10}}{6}\]

OpenStudy (jdoe0001):

$$ \sqrt{9} = \sqrt{(3)^2} = +3\\ \sqrt{9} = \sqrt{(-3)^2} = -3\\ \text{so, thus}\\ \color{blue}{\sqrt{9} = \pm 3} $$

OpenStudy (anonymous):

wat wheres the root 9 from jdoe

OpenStudy (whpalmer4):

He's just illustrating why the \(\pm\) business is there.

OpenStudy (anonymous):

so how dya wite the final answer?

OpenStudy (jdoe0001):

that is true for any value "coming out" of the square root, reason why is a \(\pm\)

OpenStudy (anonymous):

aha thnx!

OpenStudy (whpalmer4):

Well, as I wrote it is one way. Another would be \(x=\frac{1}{3}(2+\sqrt{10}), x=\frac{1}{3}(2-\sqrt{10})\)

OpenStudy (whpalmer4):

x = -0.387426, x = 1.72076 would be another

OpenStudy (whpalmer4):

I generally prefer symbolic, exact solutions (like the ones with the square root signs) to the numeric, unless a number is what is required.

OpenStudy (whpalmer4):

Do you want another one to try? I'll check your work...

OpenStudy (anonymous):

oh thnx so much but why didnt you add the 16 and 24 in the 'post 11' up

OpenStudy (whpalmer4):

Oh, I did, then I simplified it without showing the intermediary step, because I was afraid of getting cut off on the right side of the window. \[\sqrt{16+24} = \sqrt{40} = \sqrt{4*10} = \sqrt{4}*\sqrt{10} = 2\sqrt{10}\]

OpenStudy (anonymous):

OMG i have no idea how root 40 gets to 2 root 10 sorry..

OpenStudy (whpalmer4):

Even with me showing you the steps there?

OpenStudy (anonymous):

oh so it gets to 4 times 10 why cant u leave it ther

OpenStudy (whpalmer4):

How are you with exponents? Square roots are just fractional exponents, specifically\[\sqrt{x} = x^{1/2}\] We know that \[(ab)^n = a^nb^n\]so \[(ab)^{1/2} = a^{1/2}*b^{1/2}\] and so \[\sqrt{ab} = \sqrt{a}*\sqrt{b}\]

OpenStudy (whpalmer4):

You could leave it as \(\sqrt{40}\) but usually you want to simplify/reduce as much as possible.

OpenStudy (anonymous):

right thnx sososososo much!!!

OpenStudy (whpalmer4):

you're welcome!

OpenStudy (anonymous):

hey can i put in another q? how do you know when to use which formula for workin out triangles...sine,trig,cosine..

OpenStudy (whpalmer4):

that's kind of a broad question! One helpful mnemonic is SOH CAH TOA Sin = Opposite over Hypotenuse Cos = Adjacent over Hypotenuse Tan = Opposite over Adjacent "Opposite" here refers to the length of the side opposite the angle of which you are taking the sin, cos, tan; similarly "Adjacent" refers to the length of the side adjacent to the angle

OpenStudy (anonymous):

ye i know but im talkin about for which triange and to find wat eg side or angle are you supposed to use each method for?

OpenStudy (whpalmer4):

Sorry, that's not really a short answer question...

OpenStudy (anonymous):

oops ok...lets say if u know 2 angles so u use sohcahtoa 1 angle=cosine?

OpenStudy (whpalmer4):

well, SOHCAHTOA only applies if you have a right triangle. If you have arbitrary angles, then the law of sines or law of cosines might help. If you just need to find the 3rd angle, remember that the sum of the interior angles of a triangle is always 180 degrees. In general, the sum of the interior angles of an n-sided polygon is (n-2)*180, so a triangle is 180, a rectangle or trapezoid is 360, a pentagon is 540, etc.